# SOSecure: Safer Code Generation with RAG and StackOverflow Discussions

### v1

@@ Abstract.
Large Language Models (LLMs) are widely used for automated code generation. Their reliance on infrequently updated pretraining data leaves them unaware of newly discovered vulnerabilities and evolving security standards, making them prone to producing insecure code. In contrast, developer communities on Stack Overflow (SO) provide an ever-evolving repository of knowledge, where security vulnerabilities are actively discussed and addressed through collective expertise. These community-driven insights remain largely untapped by LLMs. This paper introduces SOSecure, a Retrieval-Augmented Generation (RAG) system that leverages the collective security expertise found in SO discussions to improve the security of LLM-generated code. We build a security-focused knowledge base by extracting SO answers and comments that explicitly identify vulnerabilities. Unlike common uses of RAG, SOSecure triggers after code has been generated to find discussions that identify flaws in similar code. These are used in a prompt to an LLM to consider revising the code. Evaluation across three datasets (SALLM, LLMSecEval, and LMSys) show that SOSecure achieves strong fix rates of 71.7%, 91.3%, and 96.7% respectively, compared to prompting GPT-4 without relevant discussions (49.1%, 56.5%, and 37.5%), and outperforms multiple other baselines. SOSecure operates as a language-agnostic complement to existing LLMs, without requiring retraining or fine-tuning, making it easy to deploy. Our results underscore the importance of maintaining active developer forums, which have dropped substantially in usage with LLM adoptions.
@@ 1. Introduction
LLM-powered code generation tools, such as Microsoft GitHub Copilot and OpenAI ChatGPT, have significantly improved software development efficiency (GitHub, 2025). However, these tools can inherit security flaws from their open-source training data. Programming languages and libraries evolve constantly (e.g., TensorFlow has new releases every couple of months (TensorFlow, 2025)). This evolution often involves patching vulnerabilities and replacing unsafe patterns with safe ones. A large fraction of open-source repositories, which LLMs use for pretraining, are rarely updated, leading LLMs to learn and replicate vulnerable patterns, including CWEs (common weakness enumerations) (Pearce et al., 2025). Additionally, due to capacity limitations, LLMs also often lack awareness of subtle security implications specific to particular libraries or contexts.
@@
Attackers can exploit these flaws, resulting in cyberattacks, data breaches, and degraded system performance. Despite these risks, developers frequently trust LLM-generated code without thoroughly validating its security implications, increasing the likelihood of introducing vulnerabilities into production systems (Kabir et al., 2024; Jiao et al., 2025).
@@
SO and similar developer Q&A forums present a contrasting approach to knowledge sharing. SO is a developer Q&A forum that has served as a vast repository of programming knowledge accumulated through collective expertise over more than 15 years. It encourages revising and replacing obsolete or problematic answers. Community members often highlight security concerns in comments, providing invaluable context about why certain approaches might be risky, and regularly suggesting more secure alternatives. SO answers can be updated years after they were originally posted, allowing for corrections and revisions as technologies change, making them valuable references over time. In contrast, LLMs are retrained relatively infrequently, and their answers are transient, produced on a one-off basis, typically non-deterministically, with no in-built mechanism for community validation or correction.
@@
To reduce the impact of their knowledge gaps, LLMs may use Retrieval-Augmented Generation (RAG), including from SO, before generating a response. While this might allow them to discover vulnerability-related discussions, the retrieval query is typically based on the user’s prompt, which is unlikely to elicit vulnerability-related discussions. SO also contains many older and outdated answers that may actually recommend unsafe snippets. In this work, we address these issues by proposing SOSecure: an approach focused on revising potentially vulnerable code in LLM-generated answers (and code snippets more generally) by leveraging an index of vulnerability-oriented SO discussions. SOSecure functions as a security-enhancing layer in the code generation pipeline. When a user requests code from an LLM, the code it generates is compared to relevant discussions from SOSecure’s security knowledge base, which consists of answers and comments that contain similar code patterns and explicitly mention security concerns. These retrieved discussions are then provided as additional context to the LLM along with the original code, asking if any changes would be appropriate. The LLM may then generate a revised version addressing potential security issues highlighted by the community discussions, or it may determine that no changes are necessary, if the code already follows security best practices. SOSecure works with any existing code generation LLM as a complementary security layer, leveraging the collective wisdom of the developer community to address security gaps. Our results demonstrate that this approach effectively mitigates common security vulnerabilities, and that the retrieval of relevant SO discussions is key to its success, even outperforming LLMs prompted with the specific CWE to repair.
@@
We demonstrate the generalizability of the framework by evaluating SOSecure across multiple datasets showing that our approach significantly reduces security vulnerabilities in LLM-generated code.
@@ 2.1. LLM generated code security
Several studies have assessed the security of code generated by pretrained LMs, consistently finding that all evaluated LLMs frequently produce security vulnerabilities. Recent research on ChatGPT, an instruction-tuned LM, revealed that it generates code below minimal security standards for 16 out of 21 cases and is only able to self-correct 7 cases after further prompting (Khoury et al., 2023). Despite there having been some efforts to address these issues(He and Vechev, 2023; Wang et al., 2023; Siddiq et al., 2024b, a; Chakraborty et al., 2021), security concerns in LLM-generated code remains an early-stage research topic with significant challenges.
@@ 2.2. Program Security
CWE333https://cwe.mitre.org/data/index.html is a publicly accessible classification system of common software and hardware security vulnerabilities. Each weakness type within this enumeration is assigned a unique identifier (CWE ID). Program Security involves using various tools and datasets to identify and prevent vulnerabilities. CodeQL (GitHub, 2025) and Bandit (PyCQA, 2024) represent two important static analysis approaches: CodeQL is a leading industry tool that allows custom queries for detecting security vulnerabilities across mainstream languages, while Bandit serves as a Python-specific security linter designed to identify common security issues in Python code. Both tools have proven reliable for evaluating security in LM-generated code (Siddiq et al., 2024b, a).
@@
To overcome these limitations, our work utilizes a diverse mixture of datasets collected through complementary techniques: automated data collection pipelines refined through manual inspection (SALLM (Siddiq et al., 2024b) - providing broader CWE coverage with prompts as code, LLMSecEval (Tony et al., 2023) - providing prompts as natural language), and real-world conversation datasets (LMSys (Zheng et al., 2023)) that capture authentic user interactions. This comprehensive approach allows us to evaluate security vulnerabilities across different contexts and collection methodologies.
@@ 2.4. RAG based systems
The rapidly evolving security landscape presents challenges for LLMs. In recent years, 20,000-40,000 new CVEs have been published each year (Vulnerabilities and, CVE), making it practically infeasible to include all CVE descriptions within a model’s prompt. Consequently, researchers have begun investigating strategies for adapting LLMs to evolving security vulnerabilities. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has emerged as a promising approach, enhancing generative models by incorporating relevant external information into the prompt. Du et al. (Du et al., 2024) proposed Vul-RAG, which leverages a knowledge-level RAG framework to detect vulnerabilities using existing CVE instances.
@@ 2.5. Motivating Example
When this code is generated, SOSecure analyzes it and retrieves similar code snippets from its security-aware knowledge base. Based on the similarity of code patterns, SOSecure identifies and retrieves AnswerID: 61307412444https://stackoverflow.com/a/61307412 (shown in Figure 1) . In this Stack Overflow discussion, a community member explicitly warns in a comment: “Don’t use set shell=True to run commands, it opens the program to command injection vulnerabilities.”. This comment highlights the exact security issue present in the generated code and suggests a safer alternative approach.
@@ 3. Methodology
In this study, we construct a security-aware knowledge base from SO discussions to support RAG for improving the security of LLM-generated responses. As shown in Figure 3, SOSecure starts with an existing code snippet that was previously generated by the base LLM.555It may also work on snippets from other sources, such as GitHub and Stack Overflow itself; we focus just on LLM-generated code in this paper. This code may contain security vulnerabilities. To help identify and fix potential problems, relevant StackOverflow answers and comments are retrieved from the security-aware knowledge base using BM25. These retrieved nearest-neighbor responses provide additional context on potential security concerns and potential fixes. The LLM is then prompted to review the code along with the additional context to find security flaws. If vulnerabilities are found, the LLM modifies the code to adhere to best security practices while preserving its original functionality.
@@ 3.1. StackOverflow Data Collection
We utilized the Stack Overflow data dump published in September 2024 (Exchange, 2024), which includes posts from 2008-2024. This data dump is a comprehensive collection of structured information that includes all publicly available content from the website. Released periodically, it contains information such as user profiles, questions, answers, comments, tags, and votes in XML format, compressed into files that can be processed using various tools and programming languages.
We imported these files into MySQL database tables, focusing specifically on the Posts, PostTags, and Comments tables. First, we filtered the content based on programming language using the PostTags table.
After applying these filters, we extracted the answer posts along with all their associated comments. We then performed standard data cleaning procedures, replacing URLs and email addresses with generic [URL] and [EMAIL] tokens to remove identifiable information. Using Beautiful Soup (Richardson, 2025), we removed all HTML tags except for <code> tags, which were preserved to maintain the integrity of code blocks within the posts. This ensured that programming language syntax remained intact throughout processing and analysis. Finally, we filtered for answers containing at least one code block and one comment.
@@ 3.2. Knowledge Base Construction
We implement a case-insensitive regular expression (regex) pattern to efficiently match occurrences of these keywords in Stack Overflow comments. Our secure knowledge base contains 43,338 answer posts and 38,827,772 comments for Python, and 2,000 answer posts and 1,467,317 comments for C. Our focus on comments rather than answers is deliberate, as comments often contain critical security insights from the community regarding the proposed solutions. These comments frequently highlight overlooked vulnerabilities or security considerations in otherwise functional code.
Using this approach, we construct a security-aware knowledge base where each entry is an answer with its corresponding comments- at least one of which contains security related keywords. This knowledge base effectively serves as an “antipattern” repository, capturing collective community insights by documenting instances where members identified potential security concerns.
@@ 3.3. Retrieval System
Given a code snippet, we retrieve similar answers from the security-aware knowledge base using BM25 (Robertson et al., 2009) as an information retrieval system. BM25 is a bag-of-words model that looks for how many of the query terms are present in a document. It ranks the document with the highest number of query terms, normalized by document length, at the top. We selected BM25 because previous research (Rahman et al., 2019) has demonstrated its effectiveness for code-to-code retrieval.
Using BM25 helps identify code that uses similar libraries and constructs as the reference snippet based on lexical matching. In practice, this filtering stage effectively narrows the search space to a subset of candidate code snippets that use similar libraries. We prioritize library-specific matching because security issues are often library-dependent. Although the code may be semantically similar, the use of a specific library may introduce security vulnerabilities or safety concerns that would not exist in alternative implementations.
@@ 4. Evaluation
| System | FR (%) | IR (%) | NCR (%) |
@@
4.1. Baseline Systems under test
In our study, we compare SOSecure with several baseline systems to assess its effectiveness in identifying and mitigating security vulnerabilities. The selected baselines encompass a range of approaches, each offering unique techniques that users can employ alongside simple prompting to generate secure code.
Language Model with Additional Security Prompt (GPT4+ CWE+). Building upon the previous baseline, this approach provides the language model with both the CWE name and a description of the vulnerability. By illustrating the vulnerability with a description, we aim to determine two things - whether more context about the CWE and an increased input token length enhance the model’s understanding and ability to generate appropriate fixes.

### v2

# SOSecure: The Wisdom of the Crowd for Safer AI-Generated Code

@@ Abstract.
Large Language Models (LLMs) are widely used for automated code generation. Their reliance on infrequently updated pretraining data can leave them unaware of newly discovered vulnerabilities and evolving security standards, making them prone to producing insecure code. In contrast, developer communities on Stack Overflow (SO) provide an ever-evolving repository of knowledge, where security vulnerabilities are actively discussed and addressed through collective expertise. These community-driven insights remain largely untapped by LLMs. This paper introduces SOSecure, a Retrieval-Augmented Generation (RAG) system that leverages the collective security expertise found in SO discussions to improve the security of LLM-generated code. We build a security-focused knowledge base by extracting SO answers and comments that explicitly identify vulnerabilities. Unlike common uses of RAG, SOSecure triggers after code has been generated to find discussions that identify flaws in similar code. These are used in a prompt to an LLM to consider revising the code. Evaluation across three datasets (SALLM dataset, LLMSecEval, and LMSys) shows that SOSecure achieves strong fix rates of 71.7%, 91.3%, and 96.7% respectively, compared to prompting GPT-4 without relevant discussions (49.1%, 56.5%, and 37.5%), and outperforms multiple other baselines. SOSecure operates as a language-agnostic complement to existing LLMs, without requiring retraining or fine-tuning, making it easy to deploy. Our results underscore the importance of maintaining active developer forums, which have dropped substantially in usage with LLM adoption.
@@ 1. Introduction
LLM-powered code generation tools, such as Microsoft GitHub Copilot and OpenAI ChatGPT, have significantly improved software development efficiency (GitHub, 2025b). However, these tools can inherit security flaws from their open-source training data. Programming languages and libraries evolve constantly (e.g., TensorFlow has new releases every couple of months (TensorFlow, 2025)). This evolution often involves patching vulnerabilities and replacing unsafe patterns with safe ones. A large fraction of open-source repositories, which LLMs use for pretraining, are rarely updated, leading LLMs to learn and replicate vulnerable patterns, including CWEs (Common Weakness Enumerations) (Pearce et al., 2025). Additionally, due to capacity limitations, LLMs also often lack awareness of subtle security implications specific to particular libraries or contexts.
@@
Attackers can exploit these flaws, resulting in cyberattacks, data breaches, and compromised system performance. Despite these risks, developers frequently trust LLM-generated code without thoroughly validating its security implications, increasing the likelihood of introducing vulnerabilities into production systems (Kabir et al., 2024; Jiao et al., 2025).
@@
SO and similar developer Q&A forums present a contrasting approach to knowledge sharing. SO is a developer Q&A forum that has served as a vast repository of programming knowledge accumulated through collective expertise over more than 15 years. It encourages the review and replacement of outdated or problematic answers. Community members often highlight security concerns in comments, providing invaluable context about why certain approaches might be risky, and regularly suggesting more secure alternatives. SO answers can be updated years after they were originally posted, allowing for corrections and revisions as technologies change, making them valuable references over time. In contrast, LLMs are retrained relatively infrequently, and their answers are transient, produced on a one-off basis, typically non-deterministically, with no in-built mechanism for community validation or correction.
@@
To reduce the impact of their knowledge gaps, LLMs may use Retrieval-Augmented Generation (RAG), including from SO, before generating a response. While this might allow them to discover vulnerability-related discussions, the retrieval query is typically based on the user’s prompt, which is unlikely to elicit vulnerability-related discussions. SO also contains many older and outdated answers that may actually recommend unsafe snippets. In this work, we address these issues by proposing SOSecure: an approach focused on revising potentially vulnerable code in LLM-generated answers (and code snippets more generally) leveraging an index of vulnerability-oriented SO discussions. SOSecure functions as a security-enhancing layer in the code generation pipeline. When a user requests code from an LLM, the code it generates is compared to relevant discussions from SOSecure’s security knowledge base, which consists of answers and comments that contain similar code patterns and explicitly mention security concerns. These retrieved discussions are then provided as additional context to the LLM along with the original code, asking if any changes would be appropriate. The LLM may then generate a revised version addressing potential vulnerabilities highlighted by the community discussions, or it may determine that no changes are necessary if the code already follows security best practices. SOSecure works with any existing code generation LLM as a complementary security layer, leveraging the collective wisdom of the developer community to address security gaps. Our results demonstrate that this approach effectively mitigates common security vulnerabilities and that the retrieval of relevant SO discussions is key to its success, even outperforming LLMs prompted with the specific CWE to repair.
@@
We demonstrate the generalizability of the framework by evaluating SOSecure across multiple datasets, showing that our approach significantly reduces security vulnerabilities in LLM-generated code.
@@ 2.1. LLM generated code security
Several studies have assessed the security of code generated by pretrained LMs, consistently finding that all evaluated LLMs frequently produce security vulnerabilities. Recent research on ChatGPT, an instruction-tuned LM, revealed that it generates code below minimal security standards for 16 out of 21 cases and is only able to self-correct 7 cases after further prompting (Khoury et al., 2023). Despite there having been some efforts to address these issues(He and Vechev, 2023; Wang et al., 2023; Siddiq et al., 2024b, a; Chakraborty et al., 2021), security concerns in LLM-generated code remain an early-stage research topic with significant challenges.
@@ 2.2. Program Security
CWE333https://cwe.mitre.org/data/index.html is a publicly accessible classification system of common software and hardware security vulnerabilities. Each weakness type within this enumeration is assigned a unique identifier (CWE ID). Program Security involves using various tools and datasets to identify and prevent vulnerabilities. CodeQL (GitHub, 2025a) and Bandit (PyCQA, 2024) represent two important static analysis approaches: CodeQL is a leading industry tool that allows custom queries for detecting security vulnerabilities across mainstream languages, while Bandit serves as a Python-specific security linter designed to identify common vulnerabilities in Python code. Both tools have proven reliable for evaluating security in LM-generated code (Siddiq et al., 2024b, a).
@@
To overcome these limitations, our work utilizes a diverse mixture of datasets collected through complementary techniques: automated data collection pipelines refined through manual inspection (SALLM dataset (Siddiq et al., 2024b) - providing broader CWE coverage with prompts as code, LLMSecEval (Tony et al., 2023) - providing prompts as natural language), and real-world conversation datasets (LMSys (Zheng et al., 2023)) that capture authentic user interactions. This comprehensive approach allows us to evaluate security vulnerabilities across different contexts and collection methodologies.
@@ 2.3. Crowd security discussion analysis
2.4. Security in StackOverflow Content
Previous work has shown that developer forums such as SO are a double-edged sword: While they provide accessible solutions, they also propagate insecure practices. Meng et al. (Meng et al., 2018) examined Java-related posts and found that accepted answers frequently contained insecure code, highlighting the risks of uncritical reuse. Rahman et al. (Rahman et al., 2019a) similarly identified insecure Python-related practices in SO answers, reinforcing that security issues persist across languages. More recently, Hong et al. (Hong et al., 2021) introduced DiCoS, which discovers insecure code snippets by leveraging signals from user discussions, such as comments flagging vulnerabilities. These studies underscore the need for careful filtering of SO content. Building on these insights, SOSecure explicitly selects security-related discussions and leverages community feedback to guide LLM repair.
@@ 2.5. RAG based systems
The rapidly evolving security landscape presents challenges for LLMs. In recent years, 20,000-40,000 new CVEs have been published each year (Vulnerabilities and (CVE), 2025), making it practically infeasible to include all CVE descriptions within a model’s prompt. Consequently, researchers have begun investigating strategies for adapting LLMs to evolving security vulnerabilities. Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) has emerged as a promising approach, enhancing generative models by incorporating relevant external information into the prompt. Du et al. (Du et al., 2024) proposed Vul-RAG, which leverages a knowledge-level RAG framework to detect vulnerabilities using existing CVE instances.
@@ 3. Motivating Example
When this code is generated, SOSecure analyzes it and retrieves similar code snippets from its security-aware knowledge base. Based on the similarity of code patterns, SOSecure identifies and retrieves AnswerID: 61307412444https://stackoverflow.com/a/61307412 (shown in Figure 1) . In this Stack Overflow discussion, a community member explicitly warns in a comment: “Don’t use set shell=True to run commands, it opens the program to command injection vulnerabilities.”. This comment highlights the exact vulnerability present in the generated code and suggests a safer alternative approach.
@@ 4. Methodology
In this study, we construct a security-aware knowledge base from SO discussions to support RAG to improve the security of LLM-generated responses. As shown in Figure 3, SOSecure starts with an existing code snippet that was previously generated by the base LLM.555It may also work on snippets from other sources, such as GitHub and Stack Overflow itself; we focus just on LLM-generated code in this paper. This code may contain security vulnerabilities. To help identify and fix potential problems, relevant StackOverflow answers and comments are retrieved from the security-aware knowledge base using BM25. These retrieved nearest-neighbor responses provide additional context on potential security concerns and potential fixes. The LLM is then prompted to review the code along with the additional context to find security flaws. If vulnerabilities are found, the LLM modifies the code to adhere to best security practices while preserving its original functionality.
@@ 4.1. StackOverflow Data Collection
We used the Stack Overflow data dump published in September 2024 (Exchange, 2024), which includes posts from 2008-2024. This data dump is a comprehensive collection of structured information that includes all publicly available content on the website. Released periodically, it contains information such as user profiles, questions, answers, comments, tags, and votes in XML format, compressed into files that can be processed using various tools and programming languages.
We imported these files into MySQL database tables, focusing specifically on the Posts, PostTags, and Comments tables. First, we filtered the content according to the programming language using the PostTags table.
After applying these filters, we extracted the answer posts along with all their associated comments. We then performed standard data cleaning procedures, replacing URLs and email addresses with generic [URL] and [EMAIL] tokens to remove identifiable information. Using Beautiful Soup (Richardson, 2025), we removed all HTML tags except for <code> tags, which were preserved to maintain the integrity of code blocks within posts. This ensured that the syntax of the programming language remained intact throughout the processing and analysis. Finally, we filtered for answers that contain at least one code block and one comment.
@@ 4.2. Knowledge Base Construction
We implement a case-insensitive regular expression (regex) pattern to efficiently match occurrences of these keywords in SO comments. The complete list of security-related keywords used for filtering is included in our replication package (3). Our secure knowledge base contains 43,338 answer posts and 38,827,772 comments for Python, and 2,000 answer posts and 1,467,317 comments for C. Our focus on comments rather than answers is deliberate, as comments often contain critical security insights from the community regarding the proposed solutions. These comments frequently highlight overlooked vulnerabilities or security considerations in otherwise functional code.
Using this approach, we construct a security-aware knowledge base where each entry is an answer with its corresponding comments, at least one of which contains security-related keywords. This knowledge base effectively serves as an “antipattern” repository, capturing collective community insights by documenting instances where members identified potential security concerns.
@@ 4.3. Retrieval System
Given a code snippet, we retrieve similar answers from the security-aware knowledge base using BM25 (Robertson et al., 2009) as an information retrieval system. BM25 is a bag-of-words model that looks for how many of the query terms are present in a document. It ranks the document with the highest number of query terms, normalized by document length, at the top.
We selected BM25 because previous research (Rahman et al., 2019b) has demonstrated its effectiveness for software engineering retrieval tasks, showing that it often outperforms alternatives such as VSM(Salton et al., 1975) and LSI(Landauer and Dumais, 1997) when retrieving across heterogeneous artifacts (e.g., bug reports versus source code). This observation is consistent with our own experiments. We tested dense retrieval approaches, including FAISS(Johnson et al., 2019) indexing with sentence embeddings and cosine similarity, but BM25 consistently retrieved more relevant discussions. In practice, BM25’s strength in matching exact tokens such as API names, library identifiers, and error patterns makes it a better fit for our security-aware knowledge base.
Using BM25 helps identify code that uses similar libraries and constructs as the reference snippet based on lexical matching. In practice, this filtering stage effectively narrows the search space to a subset of candidate code snippets that use similar libraries. We prioritize library-specific matching because vulnerabilities are often library-dependent. Although the code may be semantically similar, the use of a specific library may introduce security vulnerabilities or safety concerns that would not exist in alternative implementations.
@@ 5. Evaluation
5.1. Baseline Systems under test
In our study, we compare SOSecure with several baseline systems to assess its effectiveness in identifying and mitigating security vulnerabilities. The selected baselines encompass a range of approaches, from prompting-only methods (GPT4, GPT4+CWE, GPT4+CWE+), to security-focused frameworks (SALLM framework(Siddiq et al., 2024b)), to a SOTA LLM-based repair system (VulRepair(Fu et al., 2022)). Although other repair methods exist as surveyed by Zhou et al.(Zhou et al., 2025), we focus on these representative baselines because SOSecure is designed as a lightweight post-hoc retrieval layer that complements rather than replaces specialized repair models. Together, these baselines provide a balanced spectrum against which to evaluate SOSecure.
Language Model with Additional Security Prompt (GPT4+ CWE+). Building upon the previous baseline, this approach provides the language model with both the CWE name and a description of the vulnerability. By illustrating the vulnerability with a description, we aim to determine two things - whether more context about the CWE and an increased input token length enhance the model’s understanding and ability to generate appropriate fixes. Figure 4 illustrates the prompt templates used for GPT4, GPT4+CWE, GPT4+CWE+, and SOSecure.
SALLM framework. The SALLM framework provides a security-centric evaluation of LLMs. It introduces the SALLM dataset (discussed in Section 5.2), a rule-based repair component, and novel security metrics (secure@k and vulnerable@k, discussed in more detail in Section 5.3).
VulRepair. VulRepair is a SOTA LLM-based vulnerability repair system that attempts to automatically patch insecure code when prompted with vulnerability descriptions. While the underlying CodeT5 model is multilingual, the VulRepair framework and its released models have been trained and evaluated exclusively on C/C++ vulnerability datasets. For this reason, we use the LLMSecEval benchmarks with C code when comparing against VulRepair in our study.
5.2. Benchmark datasets
| System | FR | IR | NCR |
@@ 5.2. Benchmark datasets
SALLM dataset contains 100 prompts, available in both text and code formats, along with the corresponding generated code snippets covering 45 vulnerability types (CWEs). Each snippet is mapped to a CWE. We select samples that include default CodeQL QL files, resulting in a final set of 74 samples used for analysis.
LMSYS-Chat-1M consists of 1 million samples, of which 43,269 contain Python code. Among these, 31,008 samples are single-round user conversations with a single code block. To curate a high-quality dataset containing genuine vulnerabilities, we apply a two-step filtration process. First, we run the samples through two static analyzers (Bandit (PyCQA, 2024) and CodeQL) and retain only those flagged as vulnerable by both. This results in 2,809 samples. We further select samples associated with CWEs that include default CodeQL QL files, yielding a final set of 240 Python samples.
5.3. Evaluation Metrics
To determine the presence or absence of vulnerabilities in the generated code, we use CodeQL for evaluation. CodeQL is a static analysis tool designed to automatically detect vulnerabilities by executing QL queries on a database generated from the source code. We also study two programming languages, Python and C. This served as an additional motivating factor for using CodeQL, as it allows for a uniform analysis, enabling the use of a single tool to assess both languages. We compute the following security metrics to evaluate SOSecure:
Fix Rate (FR). % of vulnerabilities that were fixed by the system. A higher FR indicates better security performance.
Intro Rate (IR). % of new vulnerabilities introduced after code generation. A lower IR reflects fewer new vulnerabilities.
Fixed/Total (F/T). The raw count of vulnerabilities successfully repaired relative to the total number of vulnerable samples in the dataset.
Precision Percentage (P%). The proportion of attempted fixes that are correct. A higher P% indicates better security performance.
secure@k, vulnerable@k (Siddiq et al., 2024b) measures the security of generated code. vulnerable@k measures the probability that at least one code snippet out of k generated samples is vulnerable. For this metric, a lower score indicates better performance of the system. The secure@k metric measures the probability that all code snippets out of k samples are vulnerability-free. For this metric, a higher score indicates better performance of the system. We use both these metrics to compare the performance of SOSecure and SALLM framework.
6. Results
| Temperature | Metric | SOSecure (%) | SALLM (%) |
| 0.0 | secure@1 | 89.19 | 51.35 |
| vulnerable@1 | 10.81 | 48.65 | |
| 0.2 | secure@1 | 85.14 | 50.0 |
| vulnerable@1 | 14.86 | 50.0 | |
| 0.4 | secure@1 | 83.78 | 51.35 |
| vulnerable@1 | 16.22 | 48.65 | |
| 0.6 | secure@1 | 79.73 | 51.35 |
| vulnerable@1 | 20.27 | 48.65 | |
| 0.8 | secure@1 | 85.14 | 52.70 |
| vulnerable@1 | 14.86 | 47.30 | |
| 1.0 | secure@1 | 75.68 | 52.70 |
| vulnerable@1 | 24.32 | 47.30 |
6.1. Effectiveness of SOSecure
Table 1 presents the comparison of SOSecure with baseline approaches across three benchmark datasets. The results demonstrate that SOSecure consistently outperforms standard LLM approaches in terms of security vulnerability mitigation. On the SALLM dataset, SOSecure achieves a Fix Rate (FR) of 71.7%, significantly higher than both GPT4 (49.06%) and GPT4+CWE (58.49%). Similarly SOSecure achieves a 91.3% Fix Rate compared to 56.52% for GPT4 and 69.57% for GPT4+CWE. The most substantial performance gain is observed on the LMSys dataset, where SOSecure achieves a remarkable 96.67% Fix Rate, compared to only 37.50% for GPT4 and 45.83% for GPT4+CWE. Importantly, while improving security, SOSecure does not introduce new vulnerabilities, maintaining a 0% Introduced vulnerabilities Rate (IR) across all datasets. In contrast, GPT4+CWE introduces new vulnerabilities in 7.69% of cases on the LLMSecEval dataset. The No Change Rate (NCR) is also consistently lower for SOSecure (48.65%, 57.14%, and 3.33% across the three datasets) compared to both baseline approaches.
Beyond GPT4 variants, we also compare SOSecure against two systems designed specifically for secure code generation and repair: the SALLM framework and VulRepair.
Table 2 shows that the SALLM framework baseline, as reported in their paper (Siddiq et al., 2024b), achieves roughly 51% secure@1, whereas SOSecure consistently attains 75–89% secure@1 for varying temperatures, nearly doubling the proportion of secure generations. Vulnerable@1 also drops from about 49% with SALLM to as low as 10–15% with SOSecure.
Table 3 reports that VulRepair achieves only a 13.3% Fix Rate with a 95% No-Change Rate, leaving most vulnerabilities unaddressed. In contrast, SOSecure fixes 73.3% of vulnerabilities while introducing none (0% IR).
Overall, SOSecure outperforms all categories of baselines across datasets: it achieves higher fix rates than prompting-only techniques (GPT4, GPT4+CWE, GPT4+CWE+), stronger secure@k and lower vulnerable@k scores than the security-centric SALLM framework, and surpasses the repair system VulRepair in fix rate on the LMSys C subset. These results highlight the value of community-driven retrieval and demonstrate SOSecure’s effectiveness as a lightweight post-hoc security layer for LLM code generation.
| SOSecure | GPT4 | GPT4+CWE | VulRepair | ||||||
| CWE | Sev | F/T | P% | F/T | P% | F/T | P% | F/T | P% |
| 078 | 9.8 | 6/6 | 100 | 2/6 | 33.3 | 3/6 | 50 | 2/6 | 33.3 |
| 190 | 8.6 | 2/2 | 100 | 2/2 | 100 | 2/2 | 100 | 0/2 | 0 |
| 022 | 7.5 | 0/3 | 0 | 0/3 | 0 | 0/3 | 0 | 0/3 | 0 |
| Avg | 66.7 | 44.4 | 50 | 11.1 | |||||
| System | FR (%) | IR (%) | NCR (%) |
| SOSecure | 73.3 | 0 | 72.5 |
| GPT4 | 53.3 | 0 | 80 |
| GPT4+CWE | 60 | 0 | 77.5 |
| GPT4+CWE+ | 53.3 | 0 | 80 |
| VulRepair | 13.3 | 0 | 95 |
6.2. Performance By Vulnerability Type
In the SALLM dataset, SOSecure achieves perfect precision (100%) for CWE-918 (Server-Side Request Forgery) and CWE-089 (SQL Injection), which are critical (9.1) and high (8.8) severity vulnerabilities, respectively. For CWE-094 (Code Injection, severity 9.3), a critical vulnerability, SOSecure achieves 77.78% precision, outperforming GPT4 (55.56%) while matching GPT4+CWE. SOSecure also demonstrates superior performance for CWE-078 (OS Command Injection, severity 6.3), a medium severity vulnerability, achieving 90% precision compared to 80% for both baseline approaches.
6.3. Impact of Number of Retrieved Candidates
Figure 6 illustrates the effect of varying the number of neighbors (k) added as context on the average precision of SOSecure for LMSys data. We observe that performance improves when increasing from 1 to 3 neighbors, after which it plateaus and declines slightly beyond 5 neighbors. This suggests that excessive context may introduce noise or conflicting information, whereas too few discussions may not provide sufficient security insights. An intermediate number of neighbors therefore provides the best balance, maintaining focus on the specific vulnerability while still providing useful corrective context.